OpenStudy (anonymous):

A triangle with a base of 1/2 m has an area of 12 m^2. What is the height, in m, of the triangle?

OpenStudy (wolf1728):

triangle area = 1/2 * base * height height = area / (.5*base) height = 12 / .5*.5 height = 12 / .25 height = 48 meters

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

u r welcome